Operation Summary for May 16, 2011 - May 18, 2011 |
|
Monday May 16 The day shift began with Operations monitoring the Tevatron (TEV) store 8737 (81.64E30) and the Recycler stash (353.2E10), with the Antiproton Source (Pbar) stacking (11.2E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations established beam to the Meson FTBF T-979 experiment around 8:30 AM. Operations terminated store 8737 at 9:45 AM. Cryo system personnel began work on the wet engines at TEV sectors A3 and F3. Operators prepared CDF and D0 for accesses. TEV kicker personnel repaired the E48 kicker. Operators had the TEV in shot setup by the end of the shift. Due to an accelerated loss of helium from their target, NuMI personnel put the target helium system into flow regulation mode to prolong the target’s life, but the leak rate was still too high to keep running. NuMI will be off until another target can be prepared and installed – they estimate it will be three to four weeks before they can resume running. During the evening shift, Operations put the TEV into shot setup and established store 8741 at 5:15 PM with an average initial luminosity (AIL) of 369.13E30. Operations disabled T-979 beam at 6:04 PM. Tuesday May 17 The midnight shift began with Operations monitoring store 8741 (136.05E30) and the stash (293.3E10), with Pbar stacking (18.0E10), and with MiniBooNE taking beam. During the midnight shift, the TEV sector D2 helium roof relief leak sensor alarmed, followed immediately by a fire alarm for the D2 cryo system building. Operators responded along with the Fire Department and found the wet engine venting helium, which set off the fire alarm. The TEV quenched before operators could abort the store or dump the ramp. Cryo system personnel bypassed the D2 building cryo system and began their investigation. They reported that the D2 wet engine had a failed cylinder seal; cryo system personnel began repairs. Operators prepared the TEV for accesses into A-E sectors and Transfer Hall for necessary repairs. During the day shift, Operations reported that all TEV tunnel work had finished by 1 PM and operators began the TEV checkout. Cryo system personnel completed their D2 wet engine repairs at 1:40 PM. Operators continued the cool down of the D2 cryo system. Operators reported that the TEV sector C interlocks were dropped twice; operators searched and secured the enclosure. Operators prepared MI-65 for access. During the checkout, the F2 cryo system reported a quench, which was due to the Quench Protection Monitor (QPM) system rebooting itself. EE Support personnel did a double swap of the QPM Central Processing Unit (CPU) board with the F3 cryo system CPU board. Operators established beam to T-979 at 2:37 PM. During the evening shift, operators reported that the D2 cryo system cooldown completed at 6:06 PM and that operators had current on the TEV bus by 8:08 PM. Operators disabled the T-979 beam at 8:07 PM; this ended T-979’s present run in MTest. Operators completed the orbit checks at 11:45 PM and put the TEV into shot setup. Operations reduced the MiniBooNE beam intensity to 6.2E16 protons per hour in anticipation of an access on Wednesday, 5/18/11. Wednesday May 18 The midnight shift began with the TEV in shot setup, with Operations monitoring the stash (112.4E10), with Pbar stacking (234.7E10), and with MiniBooNE taking beam. During the midnight shift, Operations established store 8746 at 12:44 AM with an AIL of 334.8E30. Operations turned off MiniBooNE beam at 3 AM. TEV personnel began crystal collimator studies at 6 AM. The Plans for Wednesday The plans for today are to stack, stash, and store. MiniBooNE will be in access until noon. The Run Coordinator will allow various machine studies during the day. NuMI reports that their target change will take approximately a month to complete and resume running. |
|
More Information Comments and Suggestions |
| last modified 05/18/2011 email Fermilab |
| Security, Privacy, Legal | |